# Uncertainty-Aware Digital Twins: Robust Model Predictive Control using Time-Series Deep Quantile Learning

### v1

@@ Abstract
Digital Twins, virtual replicas of physical systems that enable real-time monitoring, model updates, predictions, and decision-making, present novel avenues for proactive control strategies for autonomous systems. However, achieving real-time decision-making in Digital Twins considering uncertainty necessitates an efficient uncertainty quantification (UQ) approach and optimization driven by accurate predictions of system behaviors, which remains a challenge for learning-based methods. This paper presents a simultaneous multi-step robust model predictive control (MPC) framework that incorporates real-time decision-making with uncertainty awareness for Digital Twin systems. Leveraging a multistep ahead predictor named Time-Series Dense Encoder (TiDE) as the surrogate model, this framework differs from conventional MPC models that provide only one-step ahead predictions. In contrast, TiDE can predict future states within the prediction horizon in a one-shot, significantly accelerating MPC. Furthermore, quantile regression is employed with the training of TiDE to perform flexible while computationally efficient UQ on data uncertainty. Consequently, with the deep learning quantiles, the robust MPC problem is formulated into a deterministic optimization problem and provides a safety buffer that accommodates disturbances to enhance constraint satisfaction rate. As a result, the proposed method outperforms existing robust MPC methods by providing less-conservative UQ and has demonstrated efficacy in an engineering case study involving Directed Energy Deposition (DED) additive manufacturing. This proactive while uncertainty-aware control capability positions the proposed method as a potent tool for future Digital Twin applications and real-time process control in engineering systems.
@@ 1.1 Problem Definition
The concept of Digital Twins [1, 2] has shown promising revolutions in autonomous industries such as manufacturing [3, 4, 5] and predictive maintenance [6]. It brings the idea of building bi-directional interactions between the physical system and its virtual counterpart. This enables online decision-making processes to be conducted automatically utilizing the state prediction provided by the virtual systems, and reacts proactively in response to the feedback from the physical systems [7]. One embodiment of online decision-making is via model predictive control (MPC) [8], which optimizes system performance by predicting future behavior and adjusting control inputs in real-time based on the model prediction. To account for disturbances in MPC, a family of uncertainty-aware MPC approaches has been proposed to enhance constraint satisfaction rates in the presence of anticipated uncertainty. Methods such as stochastic MPC [9] and robust MPC [10], aim to approximate uncertainty propagation through the known system dynamics, quantify the distribution of predicted states, and solve the MPC problem by explicitly incorporating uncertainty bounds with the constraints. As the accurate description of the system dynamics may be unavailable a priori, with the recent advancements in machine learning and neural networks (NN) [4], learning-based or data-driven predictive controllers [11, 12] have gained significant attention. However, although NNs can emulate the system dynamics accurately, applying NNs in uncertainty-aware MPC presents significant challenges, as quantifying and estimating uncertainty distributions can be both complex and computationally expensive. Consequently, integrating NN-based models with efficient uncertainty quantification (UQ) methods for decision-making remains an open research question critical to the advancement of Digital Twin technologies.
@@ 1.3 Data-Driven MPC
Data-driven models are essential for surrogating physics in a Digital Twin and MPC, particularly in two key scenarios: when the system’s underlying physics is overly complex or not fully understood, and when simulations are prohibitively computationally expensive or time-consuming [20]. Under these circumstances, data-driven/learning-based methods can identify the system directly using observational data. For example, neural state space models [21] can replace the system and input matrices in a state space formulation. Recursive Neural Network (RNN) and Long Short-Term Memory (LSTM) are also popular options since their structures resemble the propagation of the dynamics of the systems [22, 23, 24]. However, enabling learning-based methods with uncertainty-awareness for real-time applications is still challenging, primarily due to the computational complexity of performing UQ. Popular UQ techniques for NNs, such as ensemble methods, Bayesian NNs, Monte Carlo dropout, and bootstrapping, fall under the category of sampling-based methods [25]. While these methods can numerically approximate the distribution of NN outputs [26], their reliance on Monte Carlo sampling and multiple forward passes of NNs suffers from significant computational time, rendering them impractical for many engineering applications involving MPC.
@@
In contrast, parametric methods, which estimate the parameters of the uncertainty distribution directly, provide a computationally efficient alternative and are widely applied in learning-based MPC. For instance, Kinky Inference has been employed to learn parameters representing the bounds of system states [11]. Similarly, in [27], a Gaussian process is utilized as a discrepancy model to capture unknown system dynamics, with its predictive uncertainty serving as a probabilistic bound for nominal predictions. One noteworthy approach is quantile regression, which directly learns user-defined quantiles of the data. Unlike methods that require assumptions about the data distribution, quantile regression offers greater flexibility, making it particularly appealing for MPC application [28].
Another challenge for data-driven MPC is the computational burden of solving numerical optimizations online. Unlike linear MPC, where the closed-form solution can be analytically derived using linear quadratic regulation (LQR) [8], data-driven MPC relies on numerical iterations through black-box optimizers, requiring multiple forward passes and increasing computational efforts. To address this challenge, parallelization techniques for MPC have been proposed, focusing on accelerating online computations through improvements in both hardware and problem formulation [29]. One promising approach is the implementation of a multi-step-ahead predictor as a surrogate model. Traditional MPC uses single-step-ahead models and recursively propagates states to generate the predicted trajectory (i.e., state predictions over the entire horizon). In contrast, a multi-step-ahead predictor can provide the predicted trajectory in one-shot, significantly reducing computational overhead. A notable example is the simultaneous multi-step MPC proposed by [30], which employs a Transformer encoder as a multi-step-ahead predictor. This approach not only accelerates numerical optimization but also improves prediction accuracy by capturing temporal dependencies across multiple time steps. Moreover, compared to state-space models, multi-step-ahead predictors offer a more straightforward method for handling constraints under uncertainty. This avoids the potentially conservative sequential propagation of uncertainty seen in traditional approaches [31]. Further, although the multi-step ahead predictors have not been widely implemented in dynamical systems, the recent success in time-series forecasting have reveal promising result in efficiently quantifying data uncertainty by integrating multi-step quantile regression with time-series deep NN [32, 33]. These advancements highlight the potential of multi-step-ahead predictors to improve both computational efficiency and uncertainty handling in MPC, paving the way for the development of a robust and efficient decision-making in Digital Twins.
@@ 1.4 Research Objective
This work proposes a simultaneous multi-step robust MPC framework utilizing time-series deep neural networks and deep learning quantiles, providing a novel decision-making approach for Digital Twins of complex engineering systems. The proposed framework, depicted in Fig. 1, comprises two distinct stages. In the offline stage, noisy system data is gathered as the training data. A time-series deep neural network, named Time-Series Dense Encoder (TiDE), is employed to perform nonlinear system identification, capturing both the nominal system dynamics and the quantiles of the data uncertainty, encompassing the uncertainty of the nominal prediction. Subsequently, in the online stage, TiDE serves as the predictive model in the proposed multi-step robust MPC. The nominal prediction is utilized to assess the reference tracking performance, while the predictive quantiles are employed to guarantee constraint satisfaction. This proposed method is validated using an illustrative example and an engineering case study in additive manufacturing. The contributions of this work include:
@@
Figure 1: Proposed multi-step robust MPC framework.
@@ 2.1 MPC and Robust MPC
where $\|\mathbf{x}\|_{\mathbf{Q}}^{2}=\mathbf{x}^{\top}\mathbf{Q}\mathbf{x}$ represents the quadratic operation on the state vector $\mathbf{x}$, the weighting matrices $\mathbf{Q}\succ 0$ and $\mathbf{R}\succ 0$ are symmetric. Equation (2c) is the general representation of the dynamic equation, where $\hat{F}$ denotes the predictive model. Equations (2d) and (2e) are the constraints on states and control actions, respectively, while Equations (2f) and (2g) explicitly denote all the inequality and equality constraints if any applies.
@@
Here, $\hat{F}_{w}$ is a surrogate model trained using noisy data. $\hat{\mathbf{x}}_{k}$ is a general representation of state prediction that can either be deterministic or stochastic. Note that since the disturbance is unknown to $\hat{F}_{w}$ when making state prediction, in contrast to Equation 3, $\mathbf{w}_{k}$ is not explicitly treated as the input of $\hat{F}_{w}$.
@@ 2.2 Time Series Deep Neural Network
There are two main considerations when selecting a suitable time-series DNN for MPC: 1) The inference speed should be fast as the solving process of MPC requires several function/model evaluations, and 2) The structure of the DNN should accommodate the general format of dynamical systems as denoted in Equation (1). In this work, we select TiDE [38], illustrated in Fig. 3, as the DNN for surrogating dynamical systems, due to its forward speed, model accuracy, and the compatibility of its input structure. TiDE’s architecture, with parallelized dense layers and residual connections, ensures both computational efficiency and stable training for dynamic system modeling.
@@
$\hat{\mathbf{x}}_{k+1:k+N}=TiDE(\mathbf{x}_{k-w+1:k},\mathbf{d}_{k-w:k-1},\mathbf{u}_{k-w:k-1},\mathbf{d}_{k:k+N-1},\mathbf{u}_{k:k+N-1}|\bm{\phi}),$ (6)
where $\bm{\phi}$ is the trainable NN parameters of TiDE.
@@ 2.3 Quantile Regression
Quantile regression (or quantile loss) is a versatile statistical technique used to estimate the conditional quantiles of a response variable, such as the median or other given percentiles, based on a set of predictor variables. Unlike ordinary least-squares (OLS) regression, which focuses on modeling the mean of the response variable, quantile regression captures a broader picture by modeling the entire conditional distribution. Specifically, quantile regression is effective in quantifying aleatoric uncertainty, which arises from inherent variability in the data. One key advantage of quantile regression is its robustness to outliers, as it is less sensitive to extreme values compared to methods such as mean square error (MSE) loss. This makes it particularly useful for datasets with skewed or irregular distributions. Moreover, quantile regression does not require prior assumptions about the distribution of the data, enabling it to handle heteroscedastic uncertainty, situations where the variability of the response changes across levels of the predictors. This flexibility allows the model to adapt to complex, real-world datasets where such variability is common.

### v3

@@ Abstract
Digital Twins, virtual replicas of physical systems that enable real-time monitoring, model updates, predictions, and decision-making, present novel avenues for proactive control strategies for autonomous systems. However, achieving real-time decision-making in Digital Twins considering uncertainty necessitates an efficient uncertainty quantification (UQ) approach and optimization driven by accurate predictions of system behaviors, which remains a challenge for learning-based methods. This paper presents a simultaneous multi-step robust model predictive control (MPC) framework that incorporates real-time decision-making with uncertainty awareness for Digital Twin systems. Leveraging a multi-step-ahead predictor named Time-Series Dense Encoder (TiDE) as the surrogate model, this framework differs from conventional MPC models that provide only one-step ahead predictions. In contrast, TiDE can predict future states within the prediction horizon in one-shot, significantly accelerating MPC. Furthermore, quantile regression is employed with the training of TiDE to perform flexible while computationally efficient UQ on data uncertainty. Consequently, with the deep learning quantiles, the robust MPC problem is formulated into a deterministic optimization problem and provides a safety buffer that accommodates disturbances to enhance constraint satisfaction rate. As a result, the proposed method outperforms existing robust MPC methods by providing less-conservative UQ and has demonstrated efficacy in an engineering case study involving Directed Energy Deposition (DED) additive manufacturing. This proactive while uncertainty-aware control capability positions the proposed method as a potent tool for future Digital Twin applications and real-time process control in engineering systems.
@@ 1.1 Problem Definition
The concept of Digital Twins [1, 2] has shown promising revolutions in autonomous industries such as manufacturing [3, 4, 5] and predictive maintenance [6]. It brings the idea of building bi-directional interactions between the physical system and its virtual counterpart. This enables online decision-making processes to be conducted automatically utilizing the state prediction provided by the virtual systems, and reacts proactively in response to the feedback from the physical systems [7]. One embodiment of online decision-making for Digital Twins is via model predictive control (MPC) [8], which optimizes system performance by predicting future behavior and adjusting control inputs in real-time based on the model prediction. To account for disturbances in MPC, a family of uncertainty-aware MPC approaches has been proposed to enhance constraint satisfaction rates in the presence of anticipated uncertainty. Methods such as stochastic MPC [9] and robust MPC [10], aim to approximate uncertainty propagation through the known system dynamics, quantify the distribution of predicted states, and solve the MPC problem by explicitly incorporating uncertainty bounds with the constraints. As the accurate description of the system dynamics may be unavailable a priori, with the recent advancements in machine learning and neural networks (NN) [4], learning-based or data-driven predictive controllers [11, 12] have gained significant attention. However, although NNs can emulate the system dynamics accurately, applying NNs in uncertainty-aware MPC presents significant challenges, as quantifying and estimating uncertainty distributions can be both complex and computationally expensive. Consequently, integrating NN-based models with efficient uncertainty quantification (UQ) methods for decision-making remains an open research question critical to the advancement of Digital Twin technologies.
@@ 1.3 Data-Driven MPC
Data-driven models are essential for surrogating physics in a Digital Twin and MPC, particularly in two key scenarios: when the system’s underlying physics is overly complex or not fully understood, and when simulations are prohibitively computationally expensive or time-consuming [20]. Under these circumstances, data-driven/learning-based methods can identify the system directly using observational data. For example, neural state space models [21] can replace the system and input matrices in a state space formulation. Recursive Neural Network (RNN) and Long Short-Term Memory (LSTM) are also popular options since their structures resemble the propagation of the dynamics of the systems [22, 23, 24]. However, enabling learning-based methods with uncertainty-awareness for real-time applications is still challenging, primarily due to the computational complexity of performing UQ. Popular UQ techniques for NNs, such as ensemble methods, Bayesian NNs, Monte Carlo (MC) dropout, and bootstrapping, fall under the category of sampling-based methods [25]. While these methods can numerically approximate the distribution of NN outputs [26], their reliance on Monte Carlo sampling and multiple forward passes of NNs suffers from significant computational time, rendering them impractical for many engineering applications involving MPC.
Figure 1: Proposed multi-step robust MPC framework.
@@
In contrast, parametric methods, which estimate the parameters of the uncertainty distribution directly, provide a computationally efficient alternative and are widely applied in learning-based MPC. For instance, Kinky Inference has been employed to learn parameters representing the bounds of system states [11]. Similarly, in [27], a Gaussian process (GP) is utilized as a discrepancy model to capture unknown system dynamics, with its predictive uncertainty serving as a probabilistic bound for nominal predictions. One noteworthy approach is quantile regression, which directly learns user-defined quantiles of the data. Unlike methods that require assumptions about the data distribution, quantile regression offers greater flexibility, making it particularly appealing for MPC application [28].
Another challenge in data-driven MPC lies in the significant computational cost of solving optimization problems online. Unlike linear MPC, which benefits from efficient closed-form or QP-based solutions via linear quadratic regulation (LQR) [8], data-driven MPC relies on surrogate models, e.g. black-box functions such as neural networks, which require iterative numerical solvers and multiple forward evaluations. This becomes particularly burdensome when using one-step-ahead predictors, which must be rolled out recursively to generate the full trajectory prediction.
1.4 MPC with Multi-Step-Ahead Predictors
To alleviate the computational burden, a growing body of research has turned toward multi-step-ahead predictors, taking advantage of the recent advance in machine learning. These models generate the entire future trajectory in a single forward pass, reducing the number of function evaluations during optimization and thus significantly lowering computational cost. For example, Park [29] demonstrated that using multi-step predictors based on Transformer architectures yielded substantial improvements in runtime and prediction accuracy, especially for longer horizons. This benefit is particularly attractive in real-time MPC settings, where computational efficiency is critical.
Moreover, multi-step predictors provide structural advantages for UQ [30, 31]. Traditional robust MPC approaches, such as tube-based methods, typically require recursive uncertainty propagation, which can lead to overly conservative UQ. Recent work, such as [30], highlights that multi-step predictors allow uncertainty to be learned directly from data at the trajectory level, bypassing the need for recursive propagation and enabling simpler, more intuitive bounds. This is especially valuable in data-driven contexts, where model structure may not be fully known, and where robust decision-making must account for uncertainty without sacrificing computational tractability. Uncertainty estimation in existing methods for robust MPC with multi-step-ahead predictors [30, 32] focusing on deriving the worst-case using linear models with Set Membership identification. Although they provide intepretability and theoretical rigors in estimating erro bound, these approaches are limited in linear models and can hardly be generalized to NN-based applications.
In parallel, advances in time-series forecasting have demonstrated the effectiveness of quantile regression integrated with sequence-to-sequence deep learning models for multi-step uncertainty estimation [33, 34]. These models can capture complex temporal patterns and provide prediction intervals across time, offering a natural fit for robust decision-making that require trajectory-wise uncertainty estimates. Taken together, these trends point to multi-step-ahead prediction and direct UQ learning such as quantile regression as promising directions for improving both efficiency and robustness in data-driven MPC, particularly under real-time and uncertain environments in Digital Twin applications.
@@ 1.5 Research Objective
This work introduces a simultaneous multi-step robust MPC framework that leverages time-series deep neural networks and deep quantile learning to enable fast, uncertainty-aware decision-making in Digital Twins of complex engineering systems. While previous works have explored multi-step predictions to accelerate MPC [29] and employed quantile regression to quantify uncertainty in single-step MPC [28], this work is, to the best of our knowledge, the first to unify these two paradigms, enabling deep learning quantiles to be explicitly used as predicted bounds in multi-step robust MPC setting. By combining the computational efficiency of multi-step-ahead predictors with the flexibility and accuracy of trajectory-level uncertainty quantification, our framework opens a new direction for robust MPC that is both data-driven and scalable. As MPC is considered a popular model-based decision-making approach that can be built on existing Digital Twin frameworks, such as [4, 3, 35, 1], this proposed method pushes the limit of current MPC, providing a new option for decision-making under uncertainty in Digital Twin applications.
The proposed framework, depicted in Fig. 1, comprises two stages. In Stage 1, noisy system data is gathered as the training data. A time-series deep neural network, named Time-Series Dense Encoder (TiDE), is employed to perform nonlinear system identification, capturing both the nominal system dynamics and the quantiles of the data uncertainty, encompassing the uncertainty of the nominal prediction. Subsequently, in Stage 2, TiDE serves as the predictive model (virtual system) operating with the proposed multi-step robust MPC as the virtual-to-physical integration. The nominal prediction (median) is utilized to assess the reference tracking performance, while the predictive quantiles are employed to guarantee constraint satisfaction. This proposed method is validated using an illustrative example and an engineering case study in additive manufacturing. The contributions of this work include:
@@ 2.1 MPC and Robust MPC
where $\|\mathbf{x}\|_{\mathbf{Q}}^{2}=\mathbf{x}^{\top}\mathbf{Q}\mathbf{x}$ represents the quadratic operation on the state vector $\mathbf{x}$, the weighting matrices $\mathbf{Q}\succ 0$ and $\mathbf{R}\succ 0$ are symmetric. The MPC objective function $J$ in Eq. (2a) consists of two type of loss: $||\hat{\mathbf{x}}_{k+i+1}-\mathbf{r}_{k+i+1}||_{\mathbf{Q}}^{2}$ refers to the reference tracking error throughout the horizon using the predicted states, and $||\mathbf{u}_{k+i}||_{\mathbf{R}}^{2}$ penalizes the control efforts. The two loss terms can be balanced by user-selected $\mathbf{Q}$ and $\mathbf{R}$. Eq. (2c) is the general representation of the dynamic equation, where $\hat{F}$ denotes the predictive model. Eq. (2d) and (2e) are the constraints on states and control actions, respectively, while Equations (2f) and (2g) explicitly denote all the inequality and equality constraints if any applies.
@@
Here, $\hat{F}_{w}$ is a surrogate model trained using noisy data. $\hat{\mathbf{x}}_{k}$ is a general representation of state prediction that can either be deterministic or stochastic. Note that since the disturbance is unknown to $\hat{F}_{w}$ when making state prediction, in contrast to Eq. 3, $\mathbf{w}_{k}$ is not explicitly treated as the input of $\hat{F}_{w}$.
@@ 2.2 Time Series Deep Neural Network
There are two main considerations when selecting a suitable time-series DNN for MPC: 1) The inference speed should be fast as the solving process of MPC requires several function/model evaluations, and 2) The structure of the DNN should accommodate the general format of dynamical systems as denoted in Eq. (1). In this work, we select TiDE [40], illustrated in Fig. 3, as the DNN for surrogating dynamical systems, due to its forward speed, model accuracy, and the compatibility of its input structure. TiDE’s architecture, with parallelized dense layers and residual connections, ensures both computational efficiency and stable training for dynamic system modeling.
@@
$\hat{\mathbf{x}}_{k+1:k+N}=\text{TiDE}(\mathbf{x}_{k-w+1:k},$ $\mathbf{d}_{k-w:k-1},\mathbf{u}_{k-w:k-1},\mathbf{d}_{k:k+N-1},\mathbf{u}_{k:k+N-1}|\bm{\phi}),$ (6)
where $\bm{\phi}$ is the trainable NN parameters of TiDE.
@@ 2.3 Quantile Regression
Quantile regression (or quantile loss) [42, 43, 28] is a versatile statistical technique used to estimate the conditional quantiles of a response variable, such as the median or other given percentiles, based on a set of predictor variables. Unlike ordinary least-squares (OLS) regression, which focuses on modeling the mean of the response variable, quantile regression captures a broader picture by modeling the entire conditional distribution. Specifically, quantile regression is effective in quantifying aleatoric uncertainty, which arises from inherent variability in the data. One key advantage of quantile regression is its robustness to outliers, as it is less sensitive to extreme values compared to methods such as mean square error (MSE) loss. This makes it particularly useful for datasets with skewed or irregular distributions. Moreover, quantile regression does not require prior assumptions about the distribution of the data, enabling it to handle heteroscedastic uncertainty, situations where the variability of the response changes across levels of the predictors. This flexibility allows the model to adapt to complex, real-world datasets where such variability is common.
@@
To justify the need of quantile regression, we perform a benchmark testing against popular UQ methods, including GP, MC dropout, ensemble methods, and deep evidential regression [44]. Here, a one-dimensional benchmark problem from [44] is modified by injecting non-Gaussian noise:
$y=x^{3}+W_{exp}$ (8)
where $W_{exp}\sim\mathrm{Exp}(15)-15$ follows a zero mean Exponential distribution with rate $\lambda=15$. The reason for not assuming a normal distribution for the injected noise is that, even if the environmental disturbance is normally distributed, its propagation through a nonlinear system does not guarantee a normally distributed state. Therefore, when applying deep learning quantiles to nonlinear systems, assuming the state distribution is normal may not be the most appropriate. Aside from GP, all the NNs in each method have two hidden layers with size=64, and use ReLU() as the activation function, following the setting in [44], while the output layer may be different, depending on each case. For training the model, 1334 training samples are generated within the range of $[-5,-4]\cup[-1,4]$, and all the NN models are trained with Adam for 500 epochs with learning rate $0.1$. To evaluate the methods, a test set with size 1000 is generated using $x\in[-7,7]$ and realized using Eq. 8. Quantitative indices, including residual mean square error (RMSE), $R^{2}$, and coverage rate (num. of test samples fall within probabilistic bounds/total num. of test samples) are provided. Since the input range of training and test set are different, we not only evaluate the effecacy of learning noise distribution, but also to test the generalizability for extrapolation and interpolation at sparse data region.
Table 1: Benchmarking of UQ methods. The lower figure to each method highlights the region within the pink box.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Method Gaussian Process
Monte Carlo Dropout
Ensemble Method
UQ Type
Epistemic (CI) Epistemic (CI) Epistemic (CI)
Visualized
Bounds
[±$[\pm 1\sigma,\pm 2\sigma,\pm 3\sigma]$ [±$[\pm 1\sigma,\pm 2\sigma,\pm 3\sigma]$ [±$[\pm 1\sigma,\pm 2\sigma,\pm 3\sigma]$
Coverage Rate
for Each Bound
[10.2%, 18.2%, 29.1%] [10%, 20.8%, 36.4%] [11.1%, 23.3%, 36.3%]
RMSE/$R^{2}$ 113.1452/0.2429 56.2704/0.8127 53.9461/0.8279
Others Assume the noise is homeostatic
Dropout rate = 0.2;
MC sample=1e5
Num. of NNs: 10
[Uncaptioned image] [Uncaptioned image]
Method
Evidential Regression
Quantile Regression
UQ Type
Epistemic + Aleatoric (PI)
Aleatoric (PI)
Visualized
Bounds
[±$[\pm 1\sigma,\pm 2\sigma,\pm 3\sigma]$
$q$ = [(0.25, 0.75), (0.05, 0.95),
(0.001, 0.999)]
Coverage Rate
for Each Bound
[40.2%, 68.4%, 85.8%] [32.7%, 59.5%, 73.2%]
RMSE/$R^{2}$ 73.7554/0.6783 53.8638/0.8218
Others $\lambda=1e-4$
Each quantile is predicted
by an individual NN
As can be seen from the results in Table 1, because GP, MC dropout, ensemble method, and evidential regression all assumes the uncertainty distribution to be normal, only quantile regression can effectively capture the skewed distribution, while others failed to provide accurate uncertainty bounds. In fact, since GP, MC dropout, and ensemble method are quantifying epistemic (model) uncertainty, and are predicting confidence intervals (CI) rather than predictive intervals (PI), it is reasonable that these methods failed to capture the aleatoric (data) uncertainty, which is more critical in our application. Evidential regression, by learning the uncertainty of the estimated standard deviation, is capable of learning both epistemic and aleatoric uncertainty. However, similarly, it suffers from assuming the distribution to be normal, and provides obvious over-conservative CIs. In contrast, by using quantile regression, it can flexibly learn median response and bounds effectively regardless of the uncertainty distribution. The prediction accuracy of quantile regression also outperform other methods.
These results supports the need of quantile regression as an computationally efficient, flexible, accurate UQ methods for applications in robust MPC for nonlinear systems.
2.4 Implementation of Quantile Regression on TiDE